OpenStudy (anonymous):

Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used. All work must be shown.

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt[3]{x} -6}\]

OpenStudy (anonymous):

So far I think that I should work on the bottom half of the problem which is \[\sqrt[3]{x}-6\]

OpenStudy (anonymous):

which is \[x \frac{ -6 }{ 3 }\]

OpenStudy (anonymous):

to make it positive I \[\frac{ 1 }{ x \frac{ 6 }{ 3 } }\]

OpenStudy (anonymous):

so I think the answer is \[1/\frac{ 1 }{ x \frac{ 6 }{ 3 } }\]

OpenStudy (anonymous):

am I correct?

OpenStudy (anonymous):

@Hero can you help me please?

OpenStudy (anonymous):

@jim_thompson5910 ?

OpenStudy (anonymous):

@Directrix please?

OpenStudy (anonymous):

@ganeshie8 please?

jimthompson5910 (jim_thompson5910):

so they want you to rationalize the denominator?

OpenStudy (anonymous):

they want it to be a rational exponent form? i'm not exactly sure what that means?

jimthompson5910 (jim_thompson5910):

well why not use the idea that \[\large \sqrt[n]{x^m} = x^{m/n}\] so that means \[\large \sqrt[3]{x} = x^{1/3}\] ------------------------------------------------------- so \[\large \frac{ 1 }{ \sqrt[3]{x} -6}\] turns into \[\large \frac{ 1 }{ x^{1/3} -6}\] but I'm not sure if this is what they're looking for

OpenStudy (anonymous):

ah the -6 is an exponent inside the square root

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

so that's why I thought it was 1/1/x6/3

jimthompson5910 (jim_thompson5910):

so the original expression isn't \[\large \frac{ 1 }{ \sqrt[3]{x} -6}\] ???

OpenStudy (anonymous):

also wouldn't x^6/3 be 2x if simplified?

jimthompson5910 (jim_thompson5910):

\[\large x^{6/3} = x^{2/1} = x^2\] \[\large x^{6/3} = x^2\]

jimthompson5910 (jim_thompson5910):

x^2 is NOT the same as 2x

OpenStudy (anonymous):

no, I will take a picture of it and ohh you're right

OpenStudy (anonymous):

so would it be then 1/1/x^2?

jimthompson5910 (jim_thompson5910):

ooh i see now

OpenStudy (anonymous):

wahhhhhh

OpenStudy (anonymous):

wow

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 1 }{ \sqrt[3]{x^{-6}}}\] \[\large \frac{ 1 }{ \left(x^{-6}\right)^{1/3}}\] \[\large \frac{ 1 }{ x^{-6*1/3}}\] \[\large \frac{ 1 }{ x^{-6/3}}\] \[\large \frac{ 1 }{ x^{-2}}\] \[\large \frac{ 1 }{ \frac{1}{x^{-2}}}\] \[\large x^{2}\] ------------------------------------------------------- So \[\large \frac{ 1 }{ \sqrt[3]{x^{-6}}}\] simplifies to \[\large x^{2}\]

jimthompson5910 (jim_thompson5910):

so you can see you are very close

OpenStudy (anonymous):

but where did the 1/3 come from?

OpenStudy (anonymous):

did you separate them on purpose ?

jimthompson5910 (jim_thompson5910):

because \[\large \sqrt[3]{x} = x^{1/3}\] we know that \[\large \sqrt[3]{x^{-6}} = \left(x^{-6}\right)^{1/3}\]

OpenStudy (anonymous):

so instead of x^-6/3 you made it that way ahhhh but I get it now wow thank you very much

OpenStudy (anonymous):

do you think you can help me with one more problem?

jimthompson5910 (jim_thompson5910):

yeah there are multiple ways to get to the answer (there's probably a much easier way, but it depends on the person) and sure I can help

OpenStudy (anonymous):

thank you and here is the problem:

OpenStudy (anonymous):

One of your friends sends you a WebMail message asking you to explain how all of the following expressions have the same answer. \[\sqrt[3]{x^3}\] \[x \frac{ 1 }{ 3 } * x \frac{ 1 }{ 3 } * x \frac{ 1 }{ 3 }\] \[\frac{ 1 }{ x^1 }\] ^ the exponent 1 is negative \[\sqrt[11]{x^5 * x^4 * x^2}\] Compose a WebMail message back assisting your friend and highlight the names of the properties of exponents when you use them.

OpenStudy (anonymous):

so far I understand how the first two have the same answer but I don't understand the last two

jimthompson5910 (jim_thompson5910):

when you have a negative exponent, you flip the fraction that it applies to make the exponent positive so in general \[\large \left(\frac{a}{b}\right)^{-k} = \left(\frac{b}{a}\right)^{k}\] notice how a/b flipped to b/a and the exponent changed in sign

jimthompson5910 (jim_thompson5910):

so using this rule we can say \[\large \left(\frac{1}{x^1}\right)^{-1}\] turns into \[\large \left(\frac{x^1}{1}\right)^{1}\]

OpenStudy (anonymous):

so it would be x^1/1 whicish x^1 which is x but how is that the same as the first two

jimthompson5910 (jim_thompson5910):

well the cube root of x^3 is x this is because the cube root operation cancels out the cubing operation (kinda like how square roots undo squaring or how division undoes multiplication)

jimthompson5910 (jim_thompson5910):

example take 2 and cube it to get 2^3 = 8 then take the cube root of 8 to get 2 back again ---------------- in general, cube any number x to get x^3 then take the cube root of x^3 to get x back again

jimthompson5910 (jim_thompson5910):

so that's why \[\large \sqrt[3]{x^3} = x\]

OpenStudy (anonymous):

ahhhh your're right I looked at something wrong so the first three are the and the last one is the same because there is 11 inside the square root and outside

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

and the second one is just a different form of the first one this is because \[\large \sqrt[3]{x^3} = \left(\sqrt[3]{x}\right)^3\]

OpenStudy (anonymous):

ahh I get it all now :D and about the first question I asked you, what properties did you use to solve the problem?

OpenStudy (anonymous):

im' not very familiar with the properties yet

jimthompson5910 (jim_thompson5910):

which steps exactly are you referring to?

OpenStudy (anonymous):

the first and second ones

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

hmm I'm blanking on the proper names and only can only find pages that only list the properties themselves but they don't have names

jimthompson5910 (jim_thompson5910):

for instance the rule \[\large \sqrt[n]{x^m} = x^{m/n}\] doesn't have an official name I don't think

jimthompson5910 (jim_thompson5910):

I could try giving the rule I used for each step does that work?

OpenStudy (anonymous):

yes that would help :)

OpenStudy (anonymous):

also if you could also give me the property name of so using this rule we can say (1 x 1 ) −1 turns into (x 1 1 ) 1

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

thank you very much :)

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 1 }{ \sqrt[3]{x^{-6}}}\] \[\large \frac{ 1 }{ \left(x^{-6}\right)^{1/3}}\] \[\large \frac{ 1 }{ x^{-6*1/3}}\] \[\large \frac{ 1 }{ x^{-6/3}}\] \[\large \frac{ 1 }{ x^{-2}}\] \[\large \frac{ 1 }{ \frac{1}{x^{2}}}\] \[\large x^{2}\] Steps done Step 1: Given expression Step 2: used rule \(\large \sqrt[n]{x^m} = x^{m/n}\) Step 3: used rule \(\large \left(x^{a}\right)^{b} = x^{a*b}\) Note: this rule is may be called "The product of powers property" Step 4: Multiplication Step 5: Division Step 6: Used rule \(\large x^{-k} = \frac{1}{x^k}\) Step 7: Used rule \(\large \frac{1}{a/b} = \frac{b}{a}\)

OpenStudy (anonymous):

THANK YOU VERY MUCH! this helps a lot!

jimthompson5910 (jim_thompson5910):

I'm glad it does